Determine the maximum contribution that can be made to a Keogh plan in each of the following cases. In all instances, the individual is self-employed, and the self-employment tax reduction has already been taken.

  1. Self-employment income of $51,000.

  2. Self-employment income of $51,000 and wage income of $30,000.

  3. Self-employment income of $125,000.

  4. Self-employment income of $290,000.
Answer #1

Keogh plan is a tax deferred pension or retired plan for self employed individuals or those working in small companies where the contribution by a self employed person can be made based on 25% of self employed income or $56000 whichever is Lower.
In the given cases, all are self employment income and self employment tax reduction has already been taken. There maximum contribution for each case will be as follows.

1- When self employment income is $51000
Maximum contribution = [$51000 X (25/125) ] = $10200

2- When Self employment income is $51000. (No regard to Wage income)
Maximum contribution = [$51000 X (25/125)] = $10200

3- When Self employment income is $125000
Maximum contribution = [$125000 X (25/125)] = $25000

4- When Self employment income is $290000
Maximum contribution = [$290000 X (25/125)] = $58000 which is more than $56000, Hence contribution in this case will be $56000.